DPC manufactures medical testing strips in a continuous process environment. Direct materials are added at the beginning of the process. Conversion costs are added throughout the manufacturing process. The manufacturer has a computer-aided quality control step that rejects 1.5% of the production for defects. Requirements: Compute the numbers for normal and abnormal spoilage. Compute the equivalent units. Compute the cost per unit completed and transferred out of the process. Summarize the total costs to account for, compute the cost per equivalent units, and assign costs.

Dekalb Processing Company (DPC)
DPC manufactures medical testing strips. This is a continous Process enviroment.
Direct Material are added at the beginning of the process. Conversion costs are adding all during the manufacturing process. The manufacturing has a computer aided quality control step which
rejects 1.5% of the production for defects. Summary data for this period are as follows:
Physical Units Direct Material Conversion Costs
WIP , Beginning 10,000 250,000 100,000
    Degree of completion, beginning WIP 100% 50%
Started this period 12,000
    TOTAL UNITS TO ACCOUNT FOR:
Good units completed and transferred ?
WIP, ending inventory 2,000
   TOTAL UNITS ACCOUNTED FOR:
    Degree of completion of ending WIP 100% 60%
Total costs added this period $                                     500,000 $                       250,000
Normal spoliage as a % of good units 1.5 % of good units
Degree of completion of normal goods 100% 100%
Degree of completion of abnormal goods 100% 100%
Required:
1). Compute the units to account for, number of normal spoilage and number of abnormal spoilage.
2). Compute the Equivalent Units.
3). Compute the cost per unit and completed and transferred out of the process.
4). Compute the total costs to account for, compute cost per equivalent unit, and assign costs.
Answer #1

1. Units to account for:

Opening Units = 10000 units

+ Started this period = 12000 units

22000 units

(-) Closing Units = (2000) units

Completed units = 20000 units

Normal loss 20000*1.5%= 300 units

Units transfered to next process = 20000-300= 17000 units

There is no abnormal loss

2. Calculation of Equivalent units

Material

Opening WIP = Nil

Units started this period(12000*100%) = 12000

Normal Loss(300*100%) 300

Closing stock (2000*100%) 2000

Total units 14300

Conversion Cost

Opening WIP (10000*50%) 5000

Units started this period(12000*100%)= 12000

Normal Loss (300*100%) 300

Closing WIP(2000*60%) 1200

Total 18500 units

3. Cost per equivqlent unit

Total cost of material = 500000

Cost per unit = 500000/14300 = 34.96/ unit

Total conversion cost = 250000

Cost per unit = 250000/18500= 13.51 unit

4. Assignment of cost

Opening WIP

Material = Nil

Conversion cost = 5000*13.51= 67550

Completed units transfered

Material (10000*0+7000*34.96) = 244720

Conversion cost (5000+7000*13.51) = 162120

Total = 406840

Normal loss

3000*(34.96+13.51) = 145410

Closing WIP

Material (2000*34.96) = 69920

Conversion cost (1200*13.51) = 16212

Total 86132